### v1

# In-medium changes of nucleon cross sections tested in neutrino-induced reactions

Abstract
Historically studied in the context of heavy-ion collisions, the extent to which free nucleon-nucleon forces are modified in-medium remains undetermined by these data sets. Therefore, we investigate the impact of NN in-medium modifications on neutrino-nucleus cross section predictions using the GiBUU transport model. We find that including an in-medium lowering of the NN cross section and density dependence on excitation improves agreement with MicroBooNE neutrino-argon scattering data. This is observed for both proton and neutral pion spectra in charged-current muon neutrino and neutral-current single pion production datasets. The impact of collision broadening of the resonance is also investigated.
I Introduction
It is still an interesting problem if free nucleon-nucleon forces are modified when the nucleons are inside a nuclear environment. For example, meson exchanges could be modified and resonance excitations could be changed. Early predictions of such changes, for example those based on relativistic Brueckner-theory by Malfliet and collaborators [1] gave a significant weakening of the NN cross sections below the resonance excitations. Li and Machleidt then took these calculations to a new level of sophistication by using the Bonn meson-exchange model [2, 3]. Since then many investigations have been performed, partly for testing the Li-Machleidt cross sections and partly for extracting the nuclear medium effects from comparisons to experimental data. No clear picture has emerged from these studies, partly because in heavy-ion collisions, mostly used for these investigations of in-medium effects, one deals with a non-equilibrium situation and signals that are integrated over the time-development of the collision (see [4] and refs therein). In addition they are sensitive to the momentum-dependence of the NN interactions at momenta above the Fermi-momentum. There, however, nucleon resonance excitations with their own changes set in. Other studies have used the sensitivity of the width of the giant resonance to in-medium changes of the NN cross section [5, 6].
There is a class of reactions that has not been used to explore the NN in-medium changes. These are lepton induced reactions such as () or (). In these reactions the nuclear target is at rest and the densities are restricted to be between 0 and the nuclear saturation density with most of the target nucleons sitting at about . We have, therefore, looked at the recent neutrino-argon scattering data from MicroBooNE which provide ejected proton multiplicities, proton spectra, and neutral pion spectra. The purpose of this letter is to explore the sensitivity of these observables to possible in-medium changes.
The authors of Ref. [7] looked for possible in-medium effects in the MINERvA experiment. In this experiment, however, the energy is fairly high. This means many different processes can take place in the final state interaction phase and resonance excitations, with their own in-medium changes, prevail. This is not so in the MicroBooNE experiment [8] which we investigate in the present paper. In this experiment, because of the lower beam energies [9], only quasielastic scattering, meson exchange processes and excitation are relevant in the initial interaction (ISI) and the open channels during the final state interactions (FSI) are limited. Furthermore, the MicroBooNE liquid argon time projection chamber detector is able to track protons down to low kinetic energies with good precision, enabling spectra of these particles to be measured over a wide range of kinetic energies spanning from 10s up to 1000 MeV. This makes MicroBooNE data an ideal place to investigate the impact of NN in-medium changes, which will be more prominent for low energy protons.
II Method
This investigation is using the GiBUU theory framework and code in its 2023 version; details of the underlying theory and practical implementation are given in Refs. [10, 11]. The code is available for download from gibuu.hepforge.org.
The in-medium NN interactions for the FSI which are implemented in GiBUU are those of Li and Machleidt [2, 3]. They lower the elastic NN cross sections as a function of density. In addition the inelastic resonance excitation in NN collisions is modified by using the in-medium change explored by of Song and Ko [12]
it decreases the excitation of the resonance with increasing density. GiBUU treats this consistently in both the pion production and the pion absorption channels. In the present study it mainly affects the pion absorption through the FSI process .
Another in-medium change is due to the collisional broadening of the resonance. This plays a role mainly in the initial neutrino-nucleon interactions because secondary excitations during the FSI are very rare at the low energy of the MicroBooNE. For this collisional broadening GiBUU has implemented the density- and momentum-dependent collisional width for the obtained by Oset and Salcedo [13].
We note that we take all the in-medium changes directly from the papers quoted; there is no tuning involved. We also mention here that GiBUU does not contain any coherent processes.
III Comparison with experiment
Ref. [14] contains an extensive set of muon-neutrino charged current () cross section measurements on argon from the MicroBooNE experiment. These measurements are inclusive, but divide the channel into final states with and without protons above the detector’s 35 MeV kinetic energy tracking threshold. These are referred to as the “0p” and “Np” final states, respectively. Cross sections as a function of the proton kinematics and proton multiplicity of the final state are also presented. The work includes a number of comparisons of the MicroBooNE data to the “standard” GiBUU version, i.e. the one without any in-medium corrections. Similar to other MicroBooNE measurements [15, 16, 17], GiBUU demonstrates a consistent ability to describe the data.
Here we now enhance this comparison by showing also the results obtained with the in-medium corrections. As suggested in Ref. [18], we account for any bias induced by the regularization of the data in cross section extraction by multiplying our predictions by the matrix reported by the experiment. We also qualify our predictions’ agreement with the data via values computed using the reported covariance matrix, which includes correlations both between measurement bins and between bins corresponding to different cross section measurements.
III.1 Proton spectra
In Fig. 1 we first show the total cross sections for 0 proton and N proton (where N ) final states. In order to be counted the protons have to have a kinetic energy of > 35 MeV. Overall, the effect of the in-medium corrections is not dramatic. The 0p cross section is somewhat lowered when switching on the in-medium corrections and it is somewhat enhanced for the Np events. This behavior is a direct consequence of the in-medium lowering of the NN cross sections. The naive final state for a quasielastic interaction contains one muon and one proton. Thus, because quasielastic interactions dominate at MicroBooNE energies, the magnitude of the 0p channel is primarily driven by final state interactions in which the proton looses some of its kinetic energy and ends up either bound in the target nucleus or below the experimental tracking threshold. Protons can also get lost all together by a charge-changing pn collision. When the NN cross sections are lowered in medium there are less of these collisions which leads to less shift in strength from Np to 0p, as observed in Fig. 1.
This is also reflected in the multiplicity distribution shown in Fig. 2. As in Fig. 1, protons only count towards the multiplicity if they have a kinetic energy of > 35 MeV and the small differences between the values in the 0p bins are attributable to slightly different experimental cuts and efficiency corrections. Here strength is shifted from 0p to mostly 1p, with only small effects in the 2p and 2p bins. This is unsurprising given the dominance of quasielastic scattering at MicroBooNE energies. The shift in strength is favored by the data and allows the in-medium prediction to fall within 1 of the measurement on all multiplicity bins. The computed value reflects this improved agreement and decreases from 8.7/4 to 2.7/4 when the in-medium corrections are turned on.
In Fig. 3 we show the kinetic energy distribution for the most energetic final state proton in reactions. The first bin of this distribution includes all interactions without a final state proton and interactions with a final state proton below the 35 MeV tracking threshold. The cross section is significantly decreased for protons with energies below the threshold when the in-medium effects are turned on and the overall distribution is markedly enhanced in much better agreement with experiment; the decreases from 21.9/15 down to 6.2/15.
Finally, in Fig. 4 we show the angular distribution of the most energetic outgoing proton. The scattering angle is defined with respect to the incoming neutrino beam. The calculation using the free NN cross sections underestimates the distribution at forward angles and overestimates it at backwards angles. When the in-medium changes are turned on the forward behavior is improved. This is a direct consequence of the smaller NN cross sections which leads to less re-scattering to perpendicular and backwards angles.
III.2 spectra
We now also look for effects of in-medium changes of the NN cross sections in neutral current production of uncharged pions together with protons. For this investigation, we compare our predictions with MicroBooNE neutral current single pion production (NC) data from Ref. [22]. At MicroBooNE energies most of the pions are produced via the resonance so these cross sections are sensitive both to the properties of the resonance [13] in the ISI and to the resonance suppression proposed in Ref. [12] in the FSI. As in Ref. [14], Ref. [22] reports matrices, which we use to account for the regularization of the data, as well as a covariance matrix that includes correlations between all measurement bins, which we use to quantify agreement with the data. For these comparisons, we note that GiBUU does not simulate coherent pion production, for which there is data that indicates the process makes up around 10-20% of the total cross section at these energies [23, 24].
The momentum spectrum of uncharged pions produced in neutral-current interactions with any number of final-state protons is shown in Fig. 5. Here the influence of the in-medium corrections is significant, in particular at the peak of the distribution. This figure shows two curves obtained with in-medium changes. In the curve “GiBUU in medium wi Oset” the Li-Machleidt and Song-Ko in-medium changes are turned on and the broadening of the resonance given by the parametrization of Oset et al [13] is included. In the curve “GiBUU in medium no Oset”, we include the Li-Machleidt and Song-Ko in-medium changes, but turn off broadening.
Including the Song-Ko density dependence for excitation increases the NC cross section by reducing absorption primarily through , which is also impacted due to detailed balance. This decreases the probability that the is absorbed into the nucleus before it can decay to a pion, thereby increasing the yield. This is seen clearly in Fig. 5. The increase in strength around the peak of the distribution is favored by the data and results in improved shape agreement for the two predictions with the in-medium corrections, as is evident by their lower values. Removing the Oset collision broadening also leads to a higher peak value of the momentum distribution in even better agreement with experiment, with very little effect on the tail. This is a consequence of the decrease in the effective width, which raises the initial pion production [19]. Though providing better agreement with the data, this increase may be artificially covering the exclusion of coherent pion production. While there is some evidence that broadening is necessary to describe photo- and electro-nuclear reactions [20] the validity of the special Oset-Salcedo in-medium width remains to be seen.
In Fig. 6 we show the scattering angle of the for neutral current interactions that also produce a proton above the 35 MeV kinetic energy tracking threshold of the MicroBooNE detector. The in-medium corrections slightly improve the agreement with experiment and the overall angular distribution is described quite well. Here, the data is not sensitive to the in-medium change of the spectral function given in [13]. The predictions with and without the Oset broadening both fall well within the uncertainties of the measurement.
Finally, in Fig. 7 we give the same angular distribution for s, but now for events with 0 protons in the acceptance range. While turning on the in-medium changes again improves the agreement with experiment, there is still a clear underestimate of the cross section at forward angles () even when the in-medium changes are turned on. If we attribute this residual yield above the “GiBUU in medium wi Oset” as being due to coherent pion production, which is not contained in GiBUU, this would correspond to a total cross section of cm2/Ar for the process. Preforming the same exercise for the “GiBUU in medium no Oset” prediction yields cm2/Ar. These values equate to 24% and 16% of the measured semi-inclusive NC cross section, respectively, which is in line with measurements of coherent pion production at similar energies but on different targets [23, 24]. However, these values are higher than predictions from the NEUT [25] and GENIE [26] neutrino event generators, which estimate that the cross section for the process on argon is cm2/Ar and cm2/Ar, respectively. These event generator predictions are closer to the coherent cross section needed to bring “GiBUU in medium wi Oset” and “GiBUU in medium no Oset” up to the lower error band of the 0p NC data, which are cm2/Ar and cm2/Ar, respectively. Alternatively, this residual 0p deficit may also suggest a need for further in-medium modifications to pn cross sections which may have a larger influence on the 0p to Np ratio in NC production than in the CC channel due to the initial neutral current interaction being capable of producing a neutron. Neutral current coherent pion production is yet to be measured on argon, and a dedicated measurement of this process from the MicroBooNE experiment could shed additional light on this disagreement at forward angles.
III.3 Global Comparison
Beyond the results shown here, we also compared the GiBUU predictions with and without the in-medium modifications to the rest of the CC measurements reported in Ref. [14]. The in-medium predictions show comparable or improved agreement with the data in all cases. To quantify this, for both predictions, we compute a value across the entirety of measurements reported in Ref. [14]. Because “blockwise unfolding” [21] was employed in Ref. [14], these values account for the correlations between all measurement bins, including those corresponding to different variables. The GiBUU prediction with the free NN cross sections achieves a of 1059/704. Including the in-medium modification lowers this to 857/704. Similarly, for the three measurements we show in Figs. 2- 4, the decreases from 69.8/39 to 24.7/39, indicating a significant improvement in the description of this data when in-medium modifications of Li-Machleidt and Song-Ko are included.
We performed the analogous comparison for the rest of the NC results presented in Ref. [23], which also reports correlations between all measurement bins. Here, including the Li-Machleidt and Song-Ko in-medium changes improves agreement for all results. In particular, comparing the “GiBUU” prediction to the prediction “GiBUU in medium no Oset” prediction, the for the double-differential measurement of the angle and momentum improves from 20.3/24 to 9.3/24. The full blockwise results are less sensitive to these differences and only reduces from 43.0/78 to 31.9/78. For this comparison, we have excluded the 0p bins at low momentum and forward angles. The residual deficit in these bins, which we attribute partly to coherent pion production, becomes the primary driver of the value. This renders the test statistic insensitive to changes throughout the rest of phase space, hence their exclusion.
These comparisons paint a consistent picture across both sets of measurements; the in-medium modification of Li-Machleidt and Song-Ko improves agreement with experiment. Though the initial interaction in the NC channel differs from the CC channel in that it is dominated by resonant interactions rather than quasielastic interactions and mediated by the neutral current rather than the charged current, the FSI experience by the outgoing hadronic reaction products are identical. As such, it is unsurprising that the treatment of in-medium effects improves the description of data in both channels. Moreover, the appendix of another MicroBooNE result [27], which measured production in charged-current interactions, has also included comparisons to GiBUU with these in-medium modifications. Those comparisons are consistent with what we show here. The in-medium prediction better describes that data, particularly for the momentum, which shows analogous trends to Fig. 5.
IV Summary
The degree to which free nucleon-nucleon forces are modified within the nucleus remains an open question. Theoretical investigations suggest that there is a significant reduction of NN cross section in medium, but experimental investigations, mostly from heavy-ion collisions, has been unable to provide further insight. However, lepton induced reaction on at-rest nuclear targets are yet to be thoroughly explored. In particular, the growing body of cross section measurements being made on heavy nuclear targets by accelerator-based neutrino experiments has not been examined in this light. The well-established need for a robust description of the final state interactions that the initial reactant products experience as they exit the nucleus suggests that these measurements will be sensitive to modifications of nucleon-nucleon forces within the nucleus.
As such, we have utilized the GiBUU model to probe MicroBooNE neutrino-argon scattering data for sensitivity to in-medium modifications of NN cross sections. Using GiBUU, we implemented the lowering of the NN cross sections according to the work of Li and Machleidt. We also account for the density dependence of the excitation cross section according to the work of Song and Ko. Comparing these predictions and ones without in-medium modifications against MicroBooNE CC and NC measurements reveals that these data are quite sensitive in-medium effects. Including the in-medium lowering of the NN cross sections and density dependence of excitation better reproduces the measured proton and spectra. This is especially apparent at low proton and energies, which are regions of phase space significantly impacted by FSI. Overall, our investigation indicated that accounting for these modification within the nuclear medium is essential in obtaining a satisfactory description of the data.
Acknowledgements.
BB is supported by the Department of Energy, Office of Science, under Award No. DE-SC0007859.References
- [1] B. Ter Haar and R. Malfliet, Phys. Lett. B 172, 10-16 (1986) doi:10.1016/0370-2693(86)90207-8
- [2] G. Q. Li and R. Machleidt, Phys. Rev. C 48, 1702 (1993) doi:10.1103/PhysRevC.48.1702 [arXiv:nucl-th/9307028 [nucl-th]].
- [3] G. Q. Li and R. Machleidt, Phys. Rev. C 49, 566 (1994) doi:10.1103/PhysRevC.49.566 [arXiv:nucl-th/9308016 [nucl-th]].
- [4] M. Henri, O. Lopez, D. Durand, B. Borderie, R. Bougault, A. Chbihi, Q. Fable, J. D. Frankland, E. Galichet and D. Gruyer, et al. Phys. Rev. C 101, no.6, 064622 (2020) doi:10.1103/PhysRevC.101.064622
- [5] M. Di Toro, V. M. Kolomietz and A. B. Larionov, Phys. Rev. C 59, 3099-3108 (1999) doi:10.1103/PhysRevC.59.3099 [arXiv:nucl-th/9807070 [nucl-th]].
- [6] R. Wang, Z. Zhang, L. W. Chen, C. M. Ko and Y. G. Ma, Phys. Lett. B 807, 135532 (2020) doi:10.1016/j.physletb.2020.135532 [arXiv:2007.12011 [nucl-th]].
- [7] X. Lu and J. T. Sobczyk, Phys. Rev. C 99, no.5, 055504 (2019) doi:10.1103/PhysRevC.99.055504 [arXiv:1901.06411 [hep-ph]].
- [8] R. Acciarri et al. [MicroBooNE], JINST 12, P02017 (2017) doi:10.1088/1748-0221/12/02/P02017.
- [9] A.A. Aguilar-Arevalo et al. [MiniBooNE], Phys. Rev. D 79, 072002 (2009) doi:10.1103/PhysRevD.79.072002 [arXiv:0806.1449 [hep-ex]]
- [10] O. Buss, T. Gaitanos, K. Gallmeister, H. van Hees, M. Kaskulov, O. Lalakulich, A. B. Larionov, T. Leitner, J. Weil and U. Mosel, Phys. Rept. 512, 1-124 (2012) doi:10.1016/j.physrep.2011.12.001 [arXiv:1106.1344 [hep-ph]].
- [11] U. Mosel and K. Gallmeister, Phys. Rev. D 109, no.3, 033008 (2024) doi:10.1103/PhysRevD.109.033008 [arXiv:2308.16161 [nucl-th]].
- [12] T. Song and C. M. Ko, Phys. Rev. C 91, no.1, 014901 (2015) doi:10.1103/PhysRevC.91.014901
- [13] E. Oset and L. L. Salcedo, Nucl. Phys. A 468, 631-652 (1987) doi:10.1016/0375-9474(87)90185-0
- [14] P. Abratenko et al. [MicroBooNE], [[arXiv:2402.19216 [hep-ex]].
- [15] P. Abratenko et al. [MicroBooNE], Phys. Rev. Lett. 131, 101802 (2023) doi:10.1103/PhysRevLett.131.101802 [[arXiv:2301.03706 [hep-ex]]].
- [16] P. Abratenko et al. [MicroBooNE], Phys. Rev. Lett. 128, 151801 (2022) doi:10.1103/PhysRevLett.128.151801 [arXiv2110.14023 [hep-ex]].
- [17] P. Abratenko et al. [MicroBooNE], [arXiv2403.19574 [hep-ex]].
- [18] W. Tang, X. Li, X. Qian, H. Wei, and C. Zhang, JINST 12, P10002 (2017) doi:10.1088/1748-0221/12/10/P10002 [arXiv:1705.03568 [physics.data-an]].
- [19] O. Lalakulich and U. Mosel, Phys. Rev. C 87, no.1, 014602 (2013) doi:10.1103/PhysRevC.87.014602 [arXiv:1210.4717 [nucl-th]].
- [20] T. E. O. Ericson and W. Weise, Clarendon Press, 1988, ISBN 978-0-19-852008-5
- [21] S. Gardiner, [arXiv:2401.04065 [hep-ex]].
- [22] P. Abratenko et al. [MicroBooNE], [arXiv2404.10948 [hep-ex]].
- [23] Y. Kurimoto et al. [SciBooNE], Phys. Rev. D 81, 111102(R) (2010) doi:10.1103/PhysRevD.81.111102 [arXiv:1005.0059 [hep-ex]].
- [24] A.A. Aguilar-Arevalo et al. [MiniBooNE], Phys. Lett. B 664, 41 (2008) doi:10.1016/j.physletb.2008.05.006 [arXiv:0803.3423 [hep-ex]].
- [25] Hayato, Y., Pickering, L. Eur. Phys. J. Spec. Top. 230, 4469–4481 (2021) doi:10.1140/epjs/s11734-021-00287-7 [arXiv:2106.15809 [hep-ph]].
- [26] C. Andreopoulos, et al., Nucl. Instrum. Meth. A 614:87-104,2010 doi:10.1016/j.nima.2009.12.009 [arXiv:0905.2517 [hep-ph]].
- [27] P. Abratenko et al. [MicroBooNE], [arXiv:2404.09949 [hep-ex]].